OpenStudy (anonymous):

I don't know how to eliminate fractions such as 2x/5+3/4=9/10

sam (.sam.):

You have to use common denominators first

OpenStudy (anonymous):

I honestly have no clue what that means I need help with like everything. like step by step in clear instructions I just need an example of how to do this, I am seriously crying.

sam (.sam.):

Its simple, First look at the denominators, its 5 and 4 \[\frac{2x}{5}+\frac{3}{4}=\frac{9}{10}\] Then you would multiply each of them the opposite terms, like this \[\frac{2x \times 4}{5 \times 4}+\frac{3 \times 5}{4 \times 5}=\frac{9}{10}\] Now you get \[\frac{8x}{20}+\frac{15}{20}=\frac{9}{10}\] Now since the denominators are the same, combine them together \[\frac{8x+15}{20}=\frac{9}{10}\] Then multiply each side by 20, \[8x+15=18 \\ 8x=3 \\ x=?\]

OpenStudy (anonymous):

thank you so much, thank you thank you thank you !!!

sam (.sam.):

Haha yw :)